1/8 less than x = 3/5, x=?

Answer by unlockmath(1688)

Hello,
Let's set this (1/8 less than x = 3/5) up as:
x - 1/8 = 3/5
Multiply by 40 to get:
40x-5=24
Add 5 and divide by 40 to get:
x= 29/40
Make sense?
RJ
www.math-unlock.com